# Channel Estimation for Beyond Diagonal RIS via Tensor Decomposition

### v1

@@ VIII Conclusion and perspectives
The channel estimation problem for BD-RIS can be addressed from a tensor decomposition perspective. Recasting the received pilot signals as a block Tucker tensor model yields individual estimates of the involved channels by exploiting the multilinear structure of the received pilot signals. Decoupling the individual channels for BD-RIS has some key benefits. First, it provides improved CSI estimation accuracy over traditional LS estimation by capitalizing on the inherent (block) Kronecker structure of the combined channel. The gains are more pronounced as more transmit/receive antennas are available. Second, the channel estimates are obtained with significantly lower training overheads than the LS method.
The proposed BTKF and BTALS algorithms have interesting tradeoffs. When training overhead is not critical, BTKF is preferable due to its lower complexity, closed-form solution, and small processing delay. On the other hand, when minimizing the training overhead is crucial (especially for strongly connected BD-RIS), BTALS is the most attractive choice due to its flexibility in operating with considerably fewer training blocks than the LS estimator with no performance degradation. Moreover, the proposed BD-RIS training yields decoupled estimates of the individual channels up to a scalar factor, which is irrelevant to system optimization.

### v2

@@ VIII Conclusion and perspectives
The channel estimation problem for BD-RIS can be addressed from a tensor decomposition perspective. Recasting the received pilot signals as a block Tucker tensor model yields individual estimates of the involved channels by exploiting the multilinear structure of the received pilot signals. Decoupling the individual channels for BD-RIS has some key benefits. First, it improves CSI estimation accuracy over traditional LS estimation by capitalizing on the composite channel’s inherent (block) Kronecker structure. The gains become more pronounced as the number of transmit and receive antennas increases. Second, the channel estimates are obtained with significantly lower training overheads than the LS method. Perspectives include studying channel estimation methods in the presence of mutual coupling, which complicates the channel model and thus requires new BD-RIS designs to facilitate the training process. In this context, mutual coupling-aware channel training and performance optimization of the proposed channel estimation algorithms deserve further study. The applicability of the proposed approach to other BD-RIS architectures, beyond group-connected ones, is also interesting to explore. Also, the development of parametric channel estimation algorithms that explicitly exploit the geometrical structure of the involved channel matrices is of interest. Finally, the study of identifiability for block Tucker decomposition with a known (and structured) core tensor is still an open problem, and the derivation of necessary and sufficient conditions related to the minimum value of $K$ ensuring unique channel estimates deserves further research.